OpenStudy (anonymous):

Using complete sentences, explain how you would use the quadratic formula to solve x2 + 5x = -2. Why is the quadratic formula the best method to use?

OpenStudy (anonymous):

\[x^2+5x=-2\] Our first step would be to move over the 2 so we have 0=.... We obtain this by adding 2 to both sides \[x^2+5x+2=0\] Now our equation is in the form \[ax^2+bx+c=0\] Basically our zeroes are not integers so in that case factoring will not work and in this instance I think our answer will be even be an irrational number There is no way to factor out this equation so we our forced to use the quadratic formula. \[\frac{-b\pm \sqrt{b^2-4a*c}}{2a}\] Now a=1, b=5 and c=2 We plug in those values and get our answers